A farmer can sell 50 bushels of beans and 100 bushels of corn for $810, or he can sell 20 bushels of beans and 205 bushels of corn for $868.50. What is the selling price of beans and corn per bushel?
Answer by fractalier(6550)

Call the prices of bushels of beans and corn, b and c. Then we have
50b + 100c = 810 and
20b + 205c = 868.5
Multiply the top one by 2 and the bottom by five and subtract them and get
100b + 200c = 1620
-(100b + 1025c = 4342.5)
--------------------------
-825c = -2722.5
c = $3.30 per bushel corn
Substituting we get
50b + 100(3.30) = 810
50b + 330 = 810
50b = 480
b = $9.60 per bushel beans